# AppSupplemental Materials References

### v1

@@ 3.3.2 Treatment Assignment
In assigning treatment, we consider the timing of a mass shooting during the course of a given year. If a shooting happens in the first 6 months of a year, the observations for that year are considered as post-treatment. If the shooting happens in the last 6 months, the observations for that year are considered pre-treatment, and the following year is considered post-treatment. [^14] The same strategy is ultimately used in Texas and Georgia, though many shootings occur between legislative sessions and therefore treatment timing is less attuned to what month the shooting occurred. Additionally, our identification strategy requires both a pre- and post-shooting observation for legislators in order to assess the effects of the shooting on their respective voting behavior. To that end, we filter out observations for legislators who assume office in the same year that is considered as the first post-treatment period, and those whose last session in office is the final pre-treatment period. [^15] From our original sample of 264 individual legislators in California with at least one year of a Gun Control Score, we are left with 255 total individual legislators for observation, and of the 24 initially treated legislators, we are left with 18 for our analyses, which can be seen in Figure S.6.1. We drop 3 additional control legislators who switched parties during the period of study. See Section S.5 for additional details.
@@ 4 Results
We report the results of our staggered DiD estimation for California legislators in Figure 3 and Table 2. For each time period relative to treatment, [^16] Under this design, the first period of observation is dropped to account for multicollinearity. we estimate the average treatment effect on the treated (ATT) legislator’s Gun Control Score, and consider never treated legislators as our control group. We rely on inverse probability weighting and cluster our standard errors (bootstrapped over 5,000 iterations) at the legislator-level. This approach has a number of benefits compared to the stacked design in Cengiz et al. (2019), which we use as a robustness check for our main results and report in Section S.5.2.
Figure 3 suggests that the occurrence of a mass shooting within a legislator’s district does not cause a legislator to change their gun control positions, on average. This null finding persists when we restrict our models to only compare treated legislators to all other legislators (green points), and compare treated legislators to untreated legislators who are members of the same party (red and blue points). We report the formal ATT estimates for the first post-treatment time period, $t=0$, by each respective legislator group, in Table 2. In Table S.5.1, we report the full results of our staggered and stacked DiD designs across all three legislator groups.
@@
Note: This figure plots the ATT estimates for the period immediately prior to treatment, $t=-1$, the first post-treatment period, $t=0$, and the periods 1 and 2 years following treatment ($t=1$ and 2, respectively). The dashed vertical line represents the onset of treatment. 95% confidence intervals are based on standard errors clustered at the legislator level using inverse probability weighting. We confine the periods for which we report our estimates to only those that are most proximate to treatment, as we expect that these periods would be the most responsive to the mass shooting.
@@
Leg. x Chamber FE
Pre-Treatment Mean 0.005 0.306 -0.272
@@
  • Note: Robust standard errors clustered at the legislator level in parentheses. Gun Control Score is measured from -1 to 1, from least to most supportive of restrictive firearms legislation. Table reports the estimated effects of mass shootings occurring within a legislator’s district on a legislator’s Gun Control Score for the first post-treatment period, $t=0$. Individual legislators who switch parties throughout the duration of study are also removed.
@@
Across all three specifications, our results are null, and considering that the Gun Control Score is measured on a -1 to 1 scale, the model in Column 1 of Table 2 can be interpreted as a 1.45% decrease in a legislator’s support for restrictive firearms policies. The ATT in column 2 represents a 4.8% reduction among California Democrats and the estimate in column 3 translates to a 0.9% increase among California Republicans, with 95% confidence intervals spanning opposite directions. In sum, we find no statistically significant change in a California legislator’s annual Gun Control Score after a mass shooting occurs within the bounds of their district. In Section S.5.5, we report the results of these same model specifications but use a Bayesian IRT model to estimate positionality on gun policy as a robustness check (rather than using the Gun Control Score), and, again, find null results.
@@
Our decision to limit the periods for which we report the ATT is motivated by a concern that periods far after treatment are subject to some source of unobserved confounding that our models are not accounting for. As only certain “types” of legislators (i.e., legislators who consistently get re-elected, or those who go on to serve in another chamber, for instance) persist as observations into the periods far after treatment (or, inversely, “survive” to eventually be treated from periods long before the shooting), the variation in the Gun Control Scores for these legislators is potentially biased by increased (or decreased) homogeneity in the observations, and a smaller number of observations in general. This can also be thought of as a form of selection bias: we are losing observations in the tails due to legislators not being re-elected or being term-limited from holding office again. With that in mind, our estimates closer to treatment are better powered and estimates far from treatment are less precise. In the following section, we widen our definition of treatment, and we ask whether legislators who represent a district near a mass shooting experience changes in behavior — not limiting treatment assignment to only those legislators who represented the districts in which the shootings occurred.
@@ 4.2 Null Effects of Mass Shootings Regardless of Distance from Shooting Location
Note: We report the null effects of a mass shooting on the voting behavior of CA state legislators on firearm-related policy, with the staggered DiD models fit on all legislators, only Democratic legislators, and only Republican legislators — at varying distances of treatment assignment. For the models fit to legislators within 5 or 25 miles of a shooting, we consider legislators whose district boundaries intersect at any point within a 5 or 25 mile radii drawn from the location of a shooting as treated, respectively.
One concern with our previous approach is that we may have been overly restrictive in our treatment assignment. Mass shootings are highly salient events that presumably affect individuals in geographic areas that may be larger than the average legislative district in California (Hassell, Holbein and Baldwin, 2020; Reny et al., 2023). To identify legislators who may reasonably be treated (beyond those representing the districts in which the shootings occur) we turn to neighboring districts. In Section S.7, we detail the strategy through which we identified legislators that would be affected if we expanded our definition of treatment beyond “within district” — and we consider districts whose boundaries intersect with or are contained within 5 and 25 mile radii drawn from the location of the shooting, respectively.
As seen in Figure 4, as we expand the geographic range for treatment assignment, the number of grouped legislators declines as these multiple-treated units are removed from our analyses. [^17] Broadening our identification strategy poses some issues for the staggered DiD design that need to be resolved. The primary issue is that in widening the boundaries, some districts and legislators receive treatment at multiple time periods. To continue relying on never treated legislators as our control units, we drop any legislators who were treated by multiple shootings throughout their time in office. Even with this broadened definition of treatment, Figure 4 shows that the null finding holds. We do find some statistically significant estimates in the periods long after treatment, though for the same reasons mentioned in preceding sections, we are skeptical that these estimates are particularly reliable given (1) the relatively small number of legislators who remain in office for long periods in California; (2) the possibility that these estimates are driven by some unobserved confounding or are otherwise extraneous; and (3) these changes in voting behavior are not obviously caused by the treatment. Taken together, Figure 4 suggests that even when the effects of a mass shooting are assumed to be felt by legislators who are geographically proximate to a shooting — and not just the representatives for the district in which the shooting occurred — mass shootings do not appear to cause changes in a legislator’s positionality on gun control, regardless of party.
@@ 4.3 No Effect of Mass Shooting on Gun Control Score of Affected District’s Representatives
Our prior results estimated the effect of a mass shooting on the Gun Control Score of the individual legislators representing the district at the time of the shooting. In this section, we estimate the effect of a mass shooting on the Gun Control Score of any of the legislators representing a district in which a shooting occurred. In this sense, “treated” legislators are those who held office in the lower or upper chamber of an affected district, though they need not to have been in office at the time of a shooting. As we continue to rely on never-treated units as our control group, control units can only be legislators who never held office in a district that was previously or subsequently affected by a mass shooting.
Figure 5: ATT Estimates for Effect of Mass Shooting on Gun Control Score Demonstrate Null Results When Treatment is Assigned to District, Not Legislator, 2012-2022
Note: We report the null effects of a mass shooting on the voting behavior of CA state legislators on firearm-related policy, with the staggered DiD models fit on all legislators, only Democratic legislators, and only Republican legislators — who, at any time, represented a district that was affected by a mass shooting. Full results and additional details can be found in Section S.8.
One might think that mass shootings cause shifts in opinions on firearms policy among the electorate that lead to changes in the firearms-related policy positions of elected officials representing said district. Scholarship demonstrating the relative stability of public opinion on firearms-related issues would suggest that this is not the case (Baxter-King, 2024; Hassell, Holbein and Baldwin, 2020), and our results are in line with these findings. Figure 5 demonstrates that mass shootings have no effect, on average, on the Gun Control Score of state legislators representing affected districts.
@@ 5 Null Result Holds Across Additional States
To validate our null result for California lawmakers and discern whether the finding is generalizable in other contexts, we utilized GLC bill rankings paired with roll-call voting records from LegiScan in five additional states — Colorado, Florida, Georgia, Missouri, and Texas — to estimate the Gun Control Scores for an additional 1,707 state legislators between 2011 and 2022.181818Unfortunately, merging the LegiScan data and the GLC Bill Rankings led to some roll call votes not being included. This is an issue inherent with the LegiScan data as not all roll call votes are collected in every state in every year. We only report results for states with the most complete roll call data, and we drop years for which multiple votes are missing. The distribution of these scores over time and across states can be seen in Figure S.9.1.191919It is worth noting that not every state had data available for the entire period of study, though this ultimately has little bearing on our staggered DiD design. Additionally, the Texas and Georgia state legislatures meet biennially, so instead of having Gun Control Scores for every year, we have scores for legislators in these states once every two years. The addition of these five states brings the total number of mass shootings that we can study to 28, as seen in Table S.2.1. We score a total of 1,975 legislators across 6 states, 1,848 of whom are retained in our analyses, and 48 of whom are treated, and expect that the power and precision of our estimates should also improve.
Our estimation strategy largely follows the one outlined in Section 3.3.1, with the exception of interacting a state-fixed effect with the legislator- and chamber-fixed effect. The results of this analysis can be seen in Figure 6, in which we plot the ATT estimates at $t=$-1, 0, 1, and 2. Table 3 reports the formal estimates for $t=0$.
@@
Figure 6: Null Effects of Mass Shootings Across Six States Generally Holds, 2011-2022
Note: We report the null effects, on average, of a mass shooting on the voting behavior of state legislators in California, Colorado, Georgia, Florida, Missouri, and Texas on firearm-related policy, with the staggered DiD models fit on all legislators, only Democratic legislators, and only Republican legislators. 95% confidence intervals are based on standard errors clustered at the legislator level using inverse probability weighting. We restrict our estimates to the legislative period immediately before a mass shooting and the three immediate post-treatment sessions ($t=$1, 0, 1, and 2). We consider the effects of 28 shootings across 1,956 individual legislators, 48 of whom are considered treated.
Table 3: No Effect of Mass Shootings on Legislator’s Support for Gun Control Across 6 States, 2011-2022
Gun Control Score
(1) (2) (3)
Shooting 0.118 0.100 0.142
(0.058) (0.071) (0.092)
Legislator Group All Dems Reps
# of Legislators 1,848 795 1,057
# of Mass Shootings 28 17 14
Leg. x Chamber x State FE
Pre-Treatment Mean 0.032 0.227 -0.178
  • Note: Robust standard errors clustered at the legislator level in parentheses. Gun Control Score is measured from -1 to 1, from least to most supportive of restrictive firearms legislation. Table records the estimated effects of mass shootings occurring within a legislator’s district on a legislator’s Gun Control Score for the first post-treatment period, $t=0$. Independents are retained in the models including all legislators, but are dropped for the models estimating effects within party. Individual legislators who switch parties throughout the duration of study are also removed.
@@
Figure 6 demonstrates that, as in California alone, mass shootings do not appear to have any statistically significant effect on the Gun Control Score of state legislatures representing the districts in which the shooting occurs. Considering the effects of a mass shooting on treated legislators compared to members of their same party, we find statistically insignificant increases among Democrats (Table 3, Column 2: a 5% increase) and Republicans (Table 3, Column 3: a 7.1% increase), respectively. The estimate of the effect of a mass shooting on the Gun Control Score of all treated legislators, regardless of party, at $t=0$ is 0.118 (SE = 0.058), which translates into a statistically significant ($\alpha$ = 0.05) 5.9% increase in a legislator’s gun control positions, on average, in the period immediately following treatment. Taken together, our results in this section confirm the generalizability of the null effect that we estimate in California: mass shootings do not cause a change in a legislator’s positions on firearm-related policy, on average.
@@ 6 Do Shootings Produce Differential Effects?
To answer this question, we turn to a strategy under the staggered DiD design that estimates the ATT by treatment groups that are assigned according to the first post-treatment year for treated observations. Therefore, for any shooting in which the first post-treatment year is coded as 2019, for example, the group would be 2019. As our assignment of treatment timing considers when in the course of a year the shooting took place and is not entirely dependent upon the year of the shooting, we end up constructing 10 group cohorts for the 28 shootings across the six states we study, as seen in Table S.2.1. Figure S.10.1 presents the group-time ATT estimates for $t=0$ accordingly, and while some estimates are statistically significant, many are quite noisy and most are null. This suggests that there is variation in the effects of particular shootings within post-treatment period groups.
Is Figure S.10.1 evidence of individual shootings producing measurable effects? Not necessarily. One limitation of our design is that we are unable to estimate the effects of particular shootings due to the binning of treatment events by post-treatment period. Subsequent work may benefit from a more fine-grained analysis regarding the characteristics of mass shootings that produce differential effects on legislator’s behavior. While disaggregating our estimates by group and time reveals some suggestive patterns, our overall null finding regarding the effect of mass shootings on the Gun Control Score of state legislators holds.

### v2

@@ 3.3.2 Treatment Assignment
In assigning treatment, we consider the timing of a mass shooting during the course of a given year. If a shooting happens in the first 6 months of a year, the observations for that year are considered as post-treatment. If the shooting happens in the last 6 months, the observations for that year are considered pre-treatment, and the following year is considered post-treatment. [^13] The same strategy is ultimately used in Texas and Georgia, though most shootings occur between legislative sessions and therefore treatment timing is less attuned to what month the shooting occurred. Additionally, our identification strategy requires both a pre- and post-shooting observation for legislators in order to assess the effects of the shooting on their respective voting behavior. To that end, we filter out observations for legislators who assume office in the same year that is considered as the first post-treatment period, and those whose last session in office is the final pre-treatment period. [^14] From our original sample of 264 individual legislators in California with at least one year of a Gun Control Score, we are left with 255 total individual legislators for observation, and of the 24 initially treated legislators, we are left with 18 for our analyses, which can be seen in Figure A.7.1. We drop 3 additional control legislators in California who switched parties during the period of study. We similarly drop party-switchers in other states within our expanded analyses. See Section A.7 for additional details.
@@ 4 Results
We report the results of our stacked DiD estimation for California legislators in Figure 3 and Table 2. For each time period relative to treatment, [^15] Under this design, the last post-treatment period of observation (effectively, t or k = -1) is dropped to account for multicollinearity. we estimate the average treatment effect on the treated (ATT) legislator’s Gun Control Score, and consider never treated legislators as our control group.
@@
Note: This figure plots the ATT estimates for the periods prior to treatment ($k=-4$ to $-2$), the first post-treatment period ($k=0$), and the periods 1 to 3 years following treatment ($k=1$ to 3, respectively). The dashed vertical line represents the onset of treatment. 95% confidence intervals are based on robust standard errors clustered at the legislator-, chamber-, and year levels. We confine the periods for which we report our estimates to only those that are most proximate to treatment, as we expect that these periods would be the most responsive to the mass shooting.
Figure 3 suggests that the occurrence of a mass shooting within a legislator’s district does not cause a legislator to change their gun control positions, on average. This null finding persists across specifications comparing treated legislators to all other legislators (black circular points), and comparing treated legislators to untreated legislators who are members of the same party (red squares for Republican legislators and blue triangular points for Democrats, respectively). We also report the formal ATT estimates for the first post-treatment time period, $k=0$, by each respective legislator group, in Table 2. In Table A.5.1, we report the full results of our staggered and stacked DiD designs across all three legislator groups.
@@
Legislator FE
Chamber FE
Year FE
Pre-Treatment Mean 0.254 0.457 -0.241
@@
  • Note: Robust standard errors are clustered at the legislator-, chamber-, and year levels in parentheses. Gun Control Score is measured from -1 to 1, from least to most supportive of restrictive firearms legislation. Table reports the estimated effects of mass shootings occurring within a legislator’s district on a legislator’s Gun Control Score for the first post-treatment period, $k=0$. Pre-Treatment Mean is the average gun control score of control and treated legislators in the pre-treatment period.
@@
Across all three specifications, our results are null, and considering that the Gun Control Score is measured on a -1 to 1 scale, the coefficient in Column 1 of Table 2 can be interpreted as a 1.15% decrease in a legislator’s support for restrictive firearms policies. The ATT in column 2 represents a 4.7% reduction among California Democrats and the estimate in column 3 translates to a 1.1% increase among California Republicans, with 95% confidence intervals spanning opposite directions. In sum, we find no statistically significant change in a California legislator’s annual Gun Control Score after a mass shooting occurs within the bounds of their district. We find similarly null results in Section A.5, in which we rely upon the staggered DiD method. In Section A.6.1, we report the results of these same model specifications but use a Bayesian IRT model to estimate positionality on gun policy as a robustness check (rather than using the Gun Control Score), and, again, find null results.
@@
Our decision to limit the periods for which we report the ATT is motivated by a concern that periods far after treatment are subject to some source of unobserved confounding that our models are not accounting for. As only certain “types” of legislators (i.e., legislators who consistently get re-elected, or those who go on to serve in another chamber, for instance) persist as observations into the periods far after treatment (or, inversely, “survive” to eventually be treated from periods long before the shooting), the variation in the Gun Control Scores for these legislators is potentially biased by increased (or decreased) homogeneity in the observations, and a smaller number of observations in general. This can also be thought of as a form of selection bias: we are losing observations in the tails due to legislators not being re-elected or being term-limited from holding office again. With that in mind, our estimates closer to treatment are better powered and estimates far from treatment are less precise. This highlights some limitations that the stacked design features when compared to the staggered design that we use as a robustness check for our main results and report in Section A.5. For instance, our uncertainty estimates may be less realistic in the years far from treatment as a lack of observations of treated legislators may lead to narrower confidence intervals due to less variability in our observations. [^16] We find that the confidence intervals in the periods far removed from treatment are generally larger under the design from Callaway and Sant’Anna (2021). Instances of particularly narrow confidence intervals in periods far from treatment appear to instead be driven by attrition among our treated units. Due to California’s relatively strict term limit law (12 years in any legislative chamber), we are ultimately limited in assessing the long-running effects of a shooting as legislators tend to be term-limited from holding office before reaching, say, $k=10$ and having sufficient pre-treatment observations to be included in our analysis. In the following section, we widen our definition of treatment, and we ask whether legislators who represent a district near a mass shooting experience changes in behavior — not limiting treatment assignment to only those legislators who represented the districts in which the shootings occurred.
@@ 4.2 Null Effects of Mass Shootings Regardless of Distance from Shooting Location
Note: We report the null effects of a mass shooting on the voting behavior of CA state legislators on firearm-related policy, with the stacked DiD models fit on all legislators, only Democratic legislators, and only Republican legislators — at varying distances of treatment assignment. For the models fit to legislators within 5 or 25 miles of a shooting, we consider legislators whose district boundaries intersect at any point within a 5 or 25 mile radii drawn from the location of a shooting as treated, respectively.
One concern with our previous approach is that we may have been overly restrictive in our treatment assignment. Mass shootings are highly salient events that presumably affect individuals in geographic areas that may be larger than the average legislative district in California (Hassell, Holbein and Baldwin, 2020; Reny et al., 2023), which suggests that we ought to account for possible spillover effects from mass shootings. To identify legislators who may reasonably be treated (beyond those representing the districts in which the shootings occur) we turn to neighboring districts. In Section A.8, we detail the strategy through which we identified legislators that would be affected if we expanded our definition of treatment beyond just shootings occurring within a legislator’s district. In this section, we consider districts whose boundaries intersect with or are contained within 5 and 25 mile radii drawn from the location of the shooting, respectively.
As seen in Figure 4, as we expand the geographic range for treatment assignment, the number of total treated legislators within each model increases, and the total number of legislators decreases. As we expand the number of legislators that are treated by each shooting, we encounter the possibility of legislators being multiply-treated throughout our study period — in widening the boundaries, some districts and legislators receive treatment at multiple time periods. Following the framework of Cengiz et al. (2019), we retain multiply-treated units for their first treatment event, but remove them as treated units in subsequent shootings in order to avoid bias from additional exposures. This “first treatment only” approach yields a more robust estimation of the ATT in initial treatment periods without reducing our sample size. Even with this broadened definition of treatment, Figure 4 suggests that when the effects of a mass shooting are assumed to be felt by legislators who are geographically proximate to a shooting — and not just the representatives for the district in which the shooting occurred — mass shootings do not appear to cause changes in a legislator’s positionality on gun control, regardless of party.
@@ 4.3 No Effect of Mass Shooting on Gun Control Score of Affected District’s Representatives
Our prior results estimated the effect of a mass shooting on the Gun Control Score of the individual legislators representing the district at the time of the shooting. In this section, we estimate the effect of a mass shooting on the Gun Control Score of any of the legislators representing a district in which a shooting occurred. In this sense, “treated” legislators are those who held office in the lower or upper chamber of an affected district, though they need not have been in office at the time of a shooting. As we continue to rely on never-treated units as our control group, control units can only be legislators who never held office in a district that was previously or subsequently affected by a mass shooting.
Figure 5: ATT Estimates for Effect of Mass Shooting on Gun Control Score Demonstrate Null Results When Treatment is Assigned to District, 2012-2022
Note: We report the null effects of a mass shooting on the voting behavior of CA state legislators on firearm-related policy, with the stacked DiD models fit on all legislators, only Democratic legislators, and only Republican legislators — who, at any time, represented a district that was affected by a mass shooting. Full results and additional details can be found in Section A.9.
One might think that mass shootings cause shifts in opinions on firearms policy among the electorate that lead to changes in the firearms-related policy positions of elected officials representing said district. Scholarship demonstrating the relative stability of public opinion on firearms-related issues would suggest that this is not the case (Baxter-King, 2024; Hassell, Holbein and Baldwin, 2020), and our results are in line with these findings. Figure 5 demonstrates that mass shootings have no effect, on average, on the Gun Control Score of state legislators representing affected districts. See Table A.9.1 for the full results.
@@ 5 Null Result Holds Across Additional States
Figure 6: Null Effects of Mass Shootings Across Five Additional States Generally Holds, 2011-2022
Note: The left panel presents the coefficient estimates (ATT at k = 0) capturing the effect of mass shootings on the Gun Control Scores of legislators who experienced a shooting event in their district, relative to all other state legislators who did not (and never subsequently did). Study periods and the number of scored legislators vary by state due to data availability, and Gun Control Scores for Georgia and Texas are measured biannually. The right panel shows histograms of the distributions of Gun Control Scores for all legislator-year observations in each state for the entire study period. These scores should not be compared across states.
To validate our null result for California lawmakers and discern whether the finding is generalizable in other contexts, we utilized GLC bill rankings paired with roll-call voting records from LegiScan in five additional states — Colorado, Florida, Georgia, Missouri, and Texas — to estimate the Gun Control Scores for an additional 1,593 state legislators between 2011 and 2022.171717Unfortunately, merging the LegiScan data and the GLC Bill Rankings led to some roll call votes not being included. This is an issue inherent with the LegiScan data as not all roll call votes are collected in every state in every year. We only report results for states with the most complete roll call data, and we drop years for which multiple votes on known gun-related bills are missing. The addition of these five states brings the total number of mass shootings that we can study to 28, as seen in Table A.2.1.
Our estimation strategy for each state follows the one outlined in Section 3.3.1. The results of this analysis can be seen in Figure 6, in which we plot the ATT estimates at k = 0 for each state and the underlying distribution of the Gun Control Score across all legislator-years.
@@
Figure 6 demonstrates that, as in California alone, mass shootings do not appear to have any statistically significant effect on the Gun Control Score of state legislatures representing the districts in which the shooting occurs in Colorado, Florida, Missouri, and Texas. One exception is Georgia, though we take this estimate with a grain of salt due to fact that we are only able to estimate the effect of a single shooting on just one post-treatment period in Georgia. However, this may suggest that individual shootings produce statistically significant results that are washed away through the aggregation strategy underlying our design. We investigate this possibility further in the following section. Taken together, our results in this section confirm the generalizability of the null effect that we estimate in California: mass shootings do not cause a change in a legislator’s positions on firearm-related policy, on average.
@@ 6 Do Shootings Produce Differential Effects?
Table 3: Disaggregated ATT Estimates of Individual Shooting Incidents
Shooting ATT (t = 0) State N Legislators N Treated
2011 Seal Beach -0.095 (0.093) CA 110 2
2012 Oakland -0.222 (0.105) CA 110 2
2013 Santa Monica -0.654 (0.042) CA 74 1
2014 Alturas 0.145 (0.048) CA 115 2
2014 Isla Vista -0.025 (0.05) CA 115 2
2015 San Bernardino -0.14 (0.038) CA 116 2
2017 Rancho Tehama Reserve 0.406 (0.031) CA 112 2
2018 Bakersfield -0.003 (0.041) CA 95 1
2018 Thousand Oaks -0.007 (0.018) CA 96 2
2021 Orange -0.176 (0.021) CA 93 1
2021 San Jose 0.295 (0.021) CA 93 1
2012 Aurora 0.037 (0.071) CO 87 2
2021 Boulder -0.123 (0.034) CO 80 1
2021 Denver 0.096 (0.175) CO 97 2
2021 Atlanta 0.419 (0.095) GA 159 2
2013 Hialeah -1.109 (0.072) FL 110 1
2016 Orlando -0.003 (0.016) FL 135 1
2017 Fort Lauderdale 0.597 (0.017) FL 107 1
2017 Orlando 0.615 (0.024) FL 107 1
2018 Parkland 0.333 (0.203) FL 147 2
2019 Sebring 0.626 (0.04) FL 109 2
2020 Springfield 0.066 (0.093) MO 189 2
2015 Tennessee Colony -0.053 (0.092) TX 150 2
2016 Dallas 0.035 (0.035) TX 150 2
2017 Sutherland Springs -0.363 (0.025) TX 140 2
2018 Santa Fe -0.259 (0.071) TX 140 2
2019 El Paso -0.445 (0.121) TX 152 1
2019 Midland-Odessa 0.21 (0.051) TX 153 2
  • Note: Table reports the disaggregated ATT estimates for each shooting in the study at k = 0. Each estimate is obtained from an individual two-way fixed effects (TWFE) Difference-in-Differences regression for each individual shooting. Models incorporate legislator, chamber, and year fixed effects. The shooting year refers to the post-treatment period.
Although our primary focus is on the average effect of mass shootings, examining disaggregated (event-specific) estimates can reveal important heterogeneity — for example, suggesting when, where, and under what conditions legislators respond more or less strongly (or statistically significantly differently than control legislators) than the overall mean suggests. To investigate the effects of individual mass shootings, we disaggregate our stacked DiD and run distinct two-by-two TWFE DiD regressions for each shooting following our overall model (not subset to partisans). Our event-specific analysis reveals that the effects of mass shootings on legislators’ gun control positions are not monolithic, and our disaggregated estimates indicate that responses vary markedly by incident. Specifically, we find that some shootings are associated with significant increases in gun control support — evidenced by positive ATT estimates following the 2014 Alturas, 2017 Rancho Tehama, 2021 San Jose, 2021 Atlanta, 2019 Sebring, 2017 Orlando, 2017 Fort Lauderdale, and 2019 Midland-Odessa events — while other events correspond to decreases in support, such as those observed in 2012 Oakland, 2013 Santa Monica, 2015 San Bernardino, 2021 Orange, 2021 Boulder, 2013 Hialeah, 2017 Sutherland Springs, 2018 Santa Fe, and 2019 El Paso.
These divergent findings suggest that specific features of the shootings may drive the direction of the legislative response. Characteristics such as the number or race of victims (Markarian, 2024; Walker, Collingwood and Bunyasi, 2020), the demographic and political composition of the affected district (Hassell, Holbein and Baldwin, 2020), the intensity and framing of media coverage (Baxter-King, 2024), and the immediate political context may all shape whether a shooting mobilizes support for gun control reform or reinforces opposition. For example, we find that highly salient mass shootings such as the 2018 Parkland school shooting — that spurred nationwide protests supporting gun control (Hassell, Holbein and Baldwin, 2020) — produce large, positive effects. Similarly, shootings in which the perpetrator targeted minority groups and marginalized communities, as in the 2021 Atlanta shooting, lead to large and statistically significant changes in affected legislator’s support for gun control policies — but notably these results are in the opposite direction for the 2019 El Paso shooting, and null for the Pulse Nightclub shooting in Orlando (which is also the shooting with the most victims in our main text). On the other hand, highly salient shootings in conservative areas produce effects in the opposite direction, as is the case with the 2017 Sutherland Springs church shooting or the 2015 San Bernardino terror attack.
Subsequent work may benefit from a more fine-grained analysis regarding the characteristics of mass shootings that produce differential effects on legislator’s behavior. Importantly, while the overall average effect of mass shootings on legislative behavior may appear null when aggregated across multiple incidents, our results underscore that this null finding likely conceals considerable heterogeneity. By targeting the average treatment effect — thereby maximizing efficiency and minimizing spurious findings from multiple comparisons — our aggregation approach reliably shows that, on average, mass shootings do not alter legislators’ gun‐policy votes, while accompanying event‐specific estimates illuminate meaningful heterogeneity in changes to legislators’ Gun Control Scores around individual shooting incidents.